How long will it take for a 10-mCi (370 MBq) sample of 32P(t1/2 ¼ 14.3 days) and a 100-mCi(3.7 GBq)sampleof67Ga(t1/2 ¼ 3.2 days) topossess the same activity?

1
Expert's answer
2021-11-07T17:09:53-0500

"A_1=A_{01}2^{-\frac t{T_{1/2,~1}}},"

"A_2=A_{02}2^{-\frac t{T_{1/2,~2}}},"

"A_1=A_2,\implies"

"t=\frac{T_1T_2}{T_1-T_2}\log _2 \frac{A_{02}}{A_{01}}=13.7^h."